OpenStudy (anonymous):

12 + 42 + 72 + ... + (3n - 2)2 =n(6n^2-3n-1)/2

OpenStudy (anonymous):

@jdoe0001 Mathematical Induction

OpenStudy (anonymous):

Would like an elaborate answer or help me through what I;am supposed to do please?

OpenStudy (campbell_st):

well i think the problem can't be proven as it fails the test when n = 1 if you substitute n = 1 into the right hand side you get \[\frac{1(6 \times 1^2 - 3 \times 1 - 1)}{2}= 1\] so this means the sum of 1 term is 1 and the left hand side seems to say, the 1st term is 12... which doesn't seem to work... so it fails the 1st test for induction where n = 1 So can you please check that the information is correctly typed

OpenStudy (anonymous):

I got 1

OpenStudy (anonymous):

but it's actually 1^2+4^2 +7^2

OpenStudy (anonymous):

I just paste it and didn't look to see if it is correct sorry

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

try to add (3(n+1) - 2)^2 to the left side

OpenStudy (anonymous):

n(6n^2-3n-1)/2 + (3(n+1) - 2)^2

OpenStudy (anonymous):

see if you get (n+1)( 6(n+1)^2 - 3(n+1) - 1)/2

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

I have (3(k+1)-2^2=k+1(6(k+1)^2-3(k+1)-1/2

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

n(6n^2-3n-1)/2 + (3(n+1) - 2)^2 = (1/2) (n+1) (6 n^2+9 n+2) and (n+1)( 6(n+1)^2 - 3(n+1) - 1)/2 = (1/2) (n+1) (6 n^2+9 n+2)

OpenStudy (anonymous):

both sides are the same so, the formula is true

OpenStudy (anonymous):

How did you do that?

OpenStudy (anonymous):

wolfram alpha XD. It's just simple algebra so why do it by hand

OpenStudy (campbell_st):

ok... so if its 1^2 then you have proved it true to n = 1 now assume its true for n = k so find the sum of k terms Use the general form \[\frac{n(6n^2 - 3n - 1)}{2}\] so assume the kth term is \[S_{k} = \frac{k(6k^2 - 3k - 1)}{2}\] the last step to proving by induction is to show \[S_{k + 1} =S_{k} + a_{k + 1}\] so the k + 1 term is \[(3(k + 1) -2)^2\] so you job is to prove that right hand side is equal to the left hand side... \[\frac{(k +1)(6(k + 1)^2 - 3(k + 1) -1)}{2} = \frac{k(6k^2 -3k -1)}{2} + (3(k + 1) -2)^2\] hope this helps...